OpenStudy (anonymous):

What is the equation for a line that goes through points (-4,0) and (8, 12) A.y=1x+4 B.y=-1x+4 C.y=1x+6 D.y=1x+3 please explain what you did

jimthompson5910 (jim_thompson5910):

first you need the slope

jimthompson5910 (jim_thompson5910):

so you use this formula m = (y2 - y1)/(x2 - x1) to get the slope

jimthompson5910 (jim_thompson5910):

m = (y2 - y1)/(x2 - x1) m = (12 - 0)/(8 - (-4)) m = (12 - 0)/(8 + 4) m = ???

jimthompson5910 (jim_thompson5910):

Tell me what you get for the slope.

OpenStudy (anonymous):

I got 12/12 but I've never been good with slopes so it is probably wrong

jimthompson5910 (jim_thompson5910):

12/12 is correct which simplifies to 1

jimthompson5910 (jim_thompson5910):

so the slope is m = 1

jimthompson5910 (jim_thompson5910):

Now move onto the general form y = mx+b y = mx+b y = 1x + b ... plug in m = 1 0 = 1(-4) + b ... plug in x = -4 and y = 0 (from the first point (-4, 0)) 0 = -4 + b 0+4 = b 4 = b b = 4 --------------------------------- So we now know that m = 1 and b = 4 which means y = mx+b turns into y = 1x + 4

OpenStudy (anonymous):

so it would be A

jimthompson5910 (jim_thompson5910):

that is correct

jimthompson5910 (jim_thompson5910):

Notice how the point (8,12) is on this line y = 1x + 4 12 = 1(8) + 4 ... plug in x = 8 and y = 12 (from the point (8,12)) 12 = 8 + 4 12 = 12 which is true, so that verifies that (8,12) is on the line

OpenStudy (anonymous):

So if I were to show my work it would be written like this:

OpenStudy (anonymous):

m = (y2 - y1)/(x2 - x1) m = (12 - 0)/(8 - (-4)) m = (12 - 0)/(8 + 4) m = 1 y = mx+b y = 1x + b ... plug in m = 1 0 = 1(-4) + b ... plug in x = -4 and y = 0 (from the first point (-4, 0)) 0 = -4 + b 0+4 = b 4 = b b = 4 m = 1 and b = 4 y = 1x + 4

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

sry for some reason the notifications system on here isn't working (idk why)

OpenStudy (anonymous):

Its fine

jimthompson5910 (jim_thompson5910):

but yeah, you first show how to get the slope

jimthompson5910 (jim_thompson5910):

then you show how to use that slope, and one of the two given points, to find the y-intercept

OpenStudy (anonymous):

Okay. Thank you so much for the help :)

jimthompson5910 (jim_thompson5910):

you're welcome